# SegAnyPET: Universal Promptable Segmentation from Positron Emission Tomography Images

### v2

@@ 4.2 Experiments
Table 3: Generalization performance to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art segmentation foundation models for zero-shot interactive segmentation from PET images.
Generalization to Unseen Organs and Dataset. To validate the generalization performance of SegAnyPET, we conduct additional experiments of foundation models for promptable segmentation of training invisible organs. As shown in Tab. 2, we observe SegAnyPET also exhibits strong generalization performance and outperforms state-of-the-art segmentation foundation models. The experimental results demonstrate that by efficiently learning representations from large-scale PET images, SegAnyPET exhibits comparable performance for training invisible organs. Furthermore, we evaluate the proposed method on training invisible AutoPET-Organ dataset with significant distribution gap compared with PETS-5k dataset. As shown in Tab. 3, we observe that SegAnyPET can achieve satisfying generalization ability on unseen dataset.
Ablation Analysis. In Tab. 4, we conduct ablation experiments to evaluate the effectiveness of proposed training strategy. Three different training strategies are conducted including fine-tuning with mixed HQ and LQ, training with additional unsupervised regularization for LQ, and the proposed cross prompting confident learning strategy. Experimental results show that the proposed strategy can achieve better performance on both seen and unseen targets.
5 Discussion and Conclusion
In this work, we propose SegAnyPET, a modality-specific foundation model for universal promptable segmentation from 3D PET images. To this end, we collect and construct a large-scale PET segmentation dataset, PETS-5k, which consists of 5,731 three-dimensional whole-body PET images, encompassing over 1.3M 2D images to train the foundation model. To facilitate the model training with different annotation quality, we design a cross prompting confident learning strategy building on the foundation of promptable segmentation for noise-robust learning. Through comprehensive evaluations on both internal and external datasets, our model shows substantial capabilities in segmenting both seen and unseen targets using only one or few prompt points and robust generalization abilities to manage new data and tasks. Its performance significantly exceeds both state-of-the-art segmentation foundation models and task-specific fully supervised models.
Strategy Prompts Seen Unseen
Fine-Tuning 1 point 87.61 62.74
3 points 88.02 63.65
5 points 88.13 63.94
Consistency 1 point 89.28 67.52
3 points 89.46 70.46
5 points 89.49 70.97
CPCL 1 point 90.49 73.96
3 points 90.90 77.09
5 points 91.05 77.87
Table 4: Ablation analysis of different training strategies for zero-shot interactive segmentation from PET images.
While SegAnyPET boasts strong capabilities, there are still inherent limitations due to manual efforts for promptable segmentation, which could potentially reduce the usability of the model compared with fully automatic segmentation. We aim to incorporate semantic information together with positional prompts as additional support to enable efficient and precise auto-prompting for automatic segmentation in the future. Besides, despite the differences between existing structural medical images, these large-scale medical datasets [1] can still provide useful guidance for training in conjunction with PET images. In conclusion, as the first foundation model for PET images, we believe that SegAnyPET will advance the developments of universal segmentation and serve as a powerful pre-trained model to facilitate the applications to various downstream tasks for molecular imaging.
Supplementary Material
Dataset Split Annotation Targets Scans New Data New Label
Train LQ 5 target organs 5,631 $\checkmark$ $\checkmark$
PETS-5k Train HQ 5 target organs 40 $\checkmark$ $\checkmark$
Internal Test all 12 organs 60 $\checkmark$ $\checkmark$
AutoPET External Test tumor lesion 1,014
AutoPET-Organ External Test all 12 organs 100 $\checkmark$
A. Dataset Information
A.1. PETS-5k Dataset
A.2. AutoPET Dataset
A.3. Organ Selection and Annotation
A.4. Summary and Visualization
An overview information summary and visualization of the datasets used in our work are shown in Tab. 5 and Fig. 5.
B. Methodological Details
B.1. SegAnyPET Architecture
B.2. Implementation Training Details
B.3. 2D/3D Prompt Generation Strategy
B.4. Evaluation Metric
$DSC(G,S)=\frac{2|G\cap S|}{|G|+|S|}$
where $G$ represents the ground truth segmentation and $S$ represents the predicted segmentation. The DSC is calculated by taking twice the size of the intersection and dividing it by the sum of the sizes of the two sets. This normalization ensures that the coefficient ranges from 0 to 1, where 1 indicates perfect overlap between the ground truth and the predicted segmentation.
@@ C. Additional Experiments and Discussion
C.1. Generalization to Tumor Segmentation
@@
In addition to the internal and external evaluation on organ segmentation, we also compare SegAnyPET with other state-of-the-art segmentation foundation models for zero-shot tumor segmentation on AutoPET dataset. Table 6 presents the experimental results under different prompt settings. Contrary to the conclusions of organ segmentation, we observe that slice-by-slice segmentation of 2D foundation models outperforms 3D foundation models. A significant difference is that the target organs in our task are all continuous entities, while the whole-body tumors in AutoPET dataset are scattered multiple small targets located in various places, as shown in Fig. 8. Therefore, the 3D model cannot directly segment all these scattered tumors out using one or a few prompt points, which also indicates the limitations of our method for this task. As an important application scenario, we aim to enlarge the dataset with instance-level tumor annotation for training and evaluation of tumor lesion segmentation in the future.
Modality Registrated CT Registrated CT PET PET
Avg DSC 88.71 66.59 73.77 76.01 72.09 76.58 78.35 90.49 90.90 91.05
Table 7: Quantitative comparison different automatic and promptablt segmentation models for organ segmentation from CT and PET images.
C.2. Comparison with CT Segmentation
Since the data used in this work are whole-body PET/CT images, we conduct additional evaluations with models for automatic and promptable organ segmentation from CT images. Given that the segmentation is used to evaluate the organ metabolic intensity from PET, it is necessary to register the CT to the resolution of PET before segmentation. Through the experimental comparison between the registered CT and PET in Tab. 7, we observe the performance of segmentation from CT is inferior to that of direct segmentation from PET. Furthermore, developing a PET segmentation model can be compatible to different scenarios such as PET/MRI or CT-free PET with self-attenuation correction.

### v3

@@ B.4. Evaluation Metric
Table 5: Generalization performance to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art general-propose foundation models for interactive segmentation from PET images.
Modality Registrated CT Registrated CT PET PET
Avg DSC 88.71 66.59 73.77 76.01 72.09 76.58 78.35 90.49 90.90 91.05
Table 6: Quantitative comparison different automatic and promptable segmentation models for organ segmentation from CT and PET images.
@@ C. Additional Experiments and Discussion
C.1. Comparison with Segmentation from CT
Since the data used in this work are whole-body PET/CT images, we conduct additional evaluations with models for automatic and promptable organ segmentation from CT images. Given that the segmentation is used to evaluate the organ metabolic intensity from PET, it is necessary to register the CT to the resolution of PET before segmentation. We compare SegAnyPET with state-of-the-art CT segmentation model TotalSegmentator [36] and SAM-Med3D-turbo applied to both registered CT and PET images Through the experimental comparison between the registered CT and PET in Tab. 6, we observe that the performance of segmentation from CT is inferior to that of direct segmentation from PET. This performance gap, combined with the long-standing and significant concern of reducing radiation exposure, underscores the importance of developing a PET-only segmentation model which could be compatible to diverse scenarios, including PET/MRI or CT-free PET with self-attenuation correction.
As PET images often come with corresponding structural CT and MRI images for attenuation correction, we plan to extend SegAnyPET to a multi-modal scenario by utilizing the information from these modalities to assist in the segmentation procedure, thereby improving the overall performance of the segmentation process.
C.2. Generalization to Unseen Dataset for Organ Segmentation
We evaluated the generalization ability of SegAnyPET to unseen out-of-distribution AutoPET-Organ dataset with comparison to state-of-the-art general-propose foundation models for interactive segmentation. As shown in Tab. 5, we observe that SegAnyPET achieves satisfying generalization ability on OOD data and outperforms existing models across different prompt settings for training-visible organs. For training-invisible organs, while it surpasses SAM and MedSAM, it lags behind SAM-Med3D. One possible explanation is that SegAnyPET is trained on relatively few targets with 5 categories. In contrast, SAM-Med3D is trained on a large amount of 245 categories including multiple organs and lesions. Some unseen targets for SegAnyPET like aorta, lung lobes and prostate are visible to SAM-Med3D from other modalities. We observe when generalized to unseen targets or training-visible targets on OOD data, SegAnyPET outperforms existing models. However, when encountering both data and target shift, SegAnyPET exhibits slightly weaker performance. We aim to improve current design with techniques like using SuperVoxel to enhance training label diversity [10] in future work.
C.3. Preliminary Analysis on Tumor Segmentation
In addition to the internal and external evaluation on organ segmentation, we also conduct a preliminary evaluation SegAnyPET with other state-of-the-art segmentation foundation models for zero-shot tumor segmentation on AutoPET dataset. Table 7 presents the experimental results under different prompt settings. Contrary to the conclusions of organ segmentation, we observe that slice-by-slice segmentation of 2D foundation models outperforms 3D foundation models. A significant difference is that the target organs in our task are all continuous entities, while the whole-body tumors in AutoPET dataset are scattered multiple small targets located in various places, as shown in Fig. 9. Therefore, the 3D model cannot directly segment all these scattered tumors using one or a few prompt points. As an important application scenario, we aim to enlarge the dataset with instance-level tumor annotation for training and evaluation of tumor lesion segmentation in the future.